RC Meeting and Runplan, Monday, Sept 24

From HallCWiki
Jump to: navigation, search
  • Status
    • We had some tune beam in the hall on Saturday night. The kink in the beam pipes at the 5th pass extraction region was found on Sunday to be causing large BLM spikes. Kink was fixed today (S. Linac is back in Power Permit).
    • New 3C15 quad was found to be physically flipped (reversing field sign) and contributed to the large beam spot we saw at the target on Saturday night.
    • PSS randomly faults in the North Linac and is currently being fixed.
    • Accelerator estimates that beam will be back sometime this evening.
    • Hall C target is still setup for solid configuration only. Depending on how this evening/owl goes, we may begin to condense the target tomorrow morning.
  • Runplan
1. Get beam into Hall with reasonable beam spot size. 
2. Establish how high of current we may obtain for the moment (this is vague, but previously the goal is to get beam to us before holding out for only high current beam). 
3. BPM calibration, ion chamber calibration, carbon hole, usual stuff to establish beam...
4. The beam deflection test is still the first item on our agenda, but it is not a show stopper, and we can continue forward and go back to it if we encounter a problem or have a weird result. The beam deflection test is linked here: [1]
5. Trigger and detector commissioning...